An electric iron of 500 w is used on a 200v supply


1
Expert's answer
2021-03-18T03:35:12-0400

According to the equation:

P = I × V

where P - power, I - current, V - voltage.

From here, the current equals:

I = P / V = 500 W / 200 V = 2.5 A


Answer: 2.5 A